AnimeGhoul8863:

An expression is shown below: f(x) = 4x2 + 8x − 5 Part A: What are the x-intercepts of the graph of f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

AnimeGhoul8863:

@Hero

KyroKazimake:

After he helps can you please leave your question up. I wanna try it

AnimeGhoul8863:

sure

Hero:

The x intercepts are the values of x that occur when f(x) = 0.

AnimeGhoul8863:

ok

Hero:

So simply set f(x) = 0, then solve for x.

Hero:

By the way, for a quadratic function of the form \(f(x) = ax^2 + bx + c\) if \(a < 0\) then the vertex of the function will have a maximum value. If \(a > 0\) then the vertex will have a minimum value.

AnimeGhoul8863:

x intercept = (1 over 2 ,0) and (-5 over 2, 0)

Hero:

Cool, but you should show your work for it.

Hero:

I don't want to do the work, I just want to check it.

AnimeGhoul8863:

i know i have been showing u the full thing i just want to make sure its right before i show u the whole thing

Hero:

Yes, they are correct.

Hero:

The reason to show the steps is so that I know you actually attempted to solve it on your own as opposed to just graphing to get the points.

NarWhalGod:

TIGGS IS BACK

AnimeGhoul8863:

T_T i graphed cause it twas hard

AnimeGhoul8863:

i need help im confused with this once cause so many things give different answers

Hero:

AHA. So when you attempted to solve it, how did you set it up to solve it?

AnimeGhoul8863:

well f(x)=4x^2 + 8x - 5

Hero:

Yep, that was the given problem. But surely there is something you can do with it that I mentioned earlier.

AnimeGhoul8863:

.......................

AnimeGhoul8863:

urmmm i am really bad at the f(x) = ax^2+bx+c

Hero:

The first step was to set f(x) = 0, so you should have had: \(0 = 4x^2 + 8x - 5\) as your first step.

AnimeGhoul8863:

ok so 4x^2+8x-5 then you split them up right

Comerpickles:

You're covered Brennah? Or could I assist you?

AnimeGhoul8863:

im waiting for hero to get back and see

Hero:

Then you FACTOR them, yes. I don't know what you mean by SPLITing.

AnimeGhoul8863:

like make them into 2

AnimeGhoul8863:

(2x-1) (2x+5) would be it when its factored

AnimeGhoul8863:

@Hero

Hero:

Please show your work.

Hero:

I don't know why you are not showing your work for these. If you don't intend to show your point, I don't see any reason to post just the answer. You could get that from anywhere. A calculator of some sort. Showing your work is the only proof you can provide that you actually did the work.

AnimeGhoul8863:

(4x^2-2x) + (10x-5) =2x(2x-1)+5(2x-1) =(2x-1) (2x+5)

Hero:

Looks right. Gotta go. BBL

AnimeGhoul8863:

ok but be back soon cause i still havent been able to answer any of the question and its getting late

AnimeGhoul8863:

@Shadow

Shadow:

You have it factored, so set it to zero and solve for x

AnimeGhoul8863:

easier said than done

AnimeGhoul8863:

can u help me figure that out cause i need to have my work shown and im not gewd at that

Shadow:

\[(2x-1) (2x+5) = 0\] \[2x - 1 = 0\] \[2x + 5 = 0\] Solve for x

AnimeGhoul8863:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AnimeGhoul8863 x intercept = (1 over 2 ,0) and (-5 over 2, 0) \(\color{#0cbb34}{\text{End of Quote}}\) i had the answer i just get confused getting it

Shadow:

What do you mean?

AnimeGhoul8863:

i need to show my work so im confused how u go from 2x−1=0 2x+5=0 to 1 over 2 ,0) and (-5 over 2, 0)

AnimeGhoul8863:

my brain freezes at these moments and i can never figure it out

Shadow:

Did you get those numbers? Or did someone give them to you?

AnimeGhoul8863:

in the beginning i got them from a calculator and they hero was helping me find out how to find them by showing my work

Shadow:

\[(2x - 1)(2x + 5) = 0\] \[2x - 1 = 0\] \[2x = 1\] \[x = \frac{ 1 }{ 2}\] \[2x + 5 = 0\] \[2x = -5\] \[x = - \frac{ 5 }{ 2 }\]

AnimeGhoul8863:

OHHHHHHHH

AnimeGhoul8863:

so if thats the answer to part A lets move to part B

AnimeGhoul8863:

Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points)

Shadow:

Just graph it. The maximum is when the vertex is highest point of the graph, and the minimum is when the vertex is the lowest point of the graph.

AnimeGhoul8863:

AnimeGhoul8863:

(-1,-9)

AnimeGhoul8863:

the vertex is minimum because the graph shows it going into the negatives which make it the lowest point in the graph

AnimeGhoul8863:

@Shadow

AnimeGhoul8863:

am i correct?

Shadow:

Correct vertex and correct that it's the minimum, but not because it's "going into the negative." Just say it's because it's the lowest point.

AnimeGhoul8863:

ok and last one Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

Part C: when you graph if you use the vertex and x intercept it will give you your function

AnimeGhoul8863:

^correct or should i change

Hero:

The answer is probably lurking in your text somewhere.

AnimeGhoul8863:

what do u mean

Hero:

In your textbook or online book notes.

AnimeGhoul8863:

urrmmmmmmmmm

Hero:

What I mean is, when I was in school, even if I needed help with something, I didn't have access to the help I needed. Had to figure it out on my own. Our role isn't to help you with every single question.

AnimeGhoul8863:

its to help me with the questions i ask but ok ill just try to do it myself even tho this is the reason im here thx for the help u gave